# Multi-scale analysis of flow over heterogeneous urban environments

### v1

@@ Abstract
A multi-scale planar-averaging framework for urban areas is developed which enables efficient computation of coarse-grained quantities. We apply the multi-scale framework to a large-eddy simulation of an idealised heterogeneous urban environment of 512 buildings based on a typical London height distribution. We observe that for this geometry, the characteristic urban lengthscale $\ell\approx 50$ m, which is the averaging lengthscale $L$ at which as much variance in the flow is resolved as is unresolved. For $L>400$ m, the statistics become approximately homogeneous, suggesting that non-building-resolving numerical weather prediction (NWP) models can be applied without modification at resolutions of 400 m and above for the case under consideration. We derive the multi-scale plane- and Reynolds-averaged momentum equation and show that for neutral cases, NWP models require parameterisation of the distributed drag and the unresolved turbulence and dispersive stress. We show that the universal drag parameterisation from Sutzl et al., 2020, Bound-Layer Meteorol., 178: 225–248 holds reasonably well for resolutions $L$ above 200 m. Below this value the problem becomes inhomogeneous and the parameterisation works less well. The unresolved stresses are well represented by a $k-\omega$ closure with a value of $\omega=0.4\,\textnormal{s}^{-1}$. However, an even more accurate closure can be derived from the Sutzl drag parameterisation that does not require further turbulence information.
Keywords:
Drag parameterisation Heterogeneity Multi-scale analysis Urban canopy Large-eddy simulation
@@ 1 Introduction
The resolution of Numerical Weather Prediction (NWP) models has advanced significantly over recent decades, transitioning from synoptic-scale $O$(10 km) grids (Bryan et al. 2003) to convection-permitting $O$(1 km) systems (Baldauf et al. 2011; Tang et al. 2013). This progress, driven by computational advancements, enables explicit resolution of critical atmospheric processes such as deep convection and orographic flows, reducing reliance on parameterizations (Lean et al. 2024). Kilometre-scale models are now operational for weather forecasting, regional climate downscaling (Prein et al. 2015; Belušić et al. 2020), and air quality assessments. However, a persistent limitation across NWP frameworks is the unresolved representation of urban morphology-buildings, vegetation, and surface heterogeneities that remain subgrid-scale features, necessitating parameterized approximations.
@@
The urban canopy region is at the lowest level of the atmospheric boundary layer and is characterised by strong heterogeneity and complex interactions between the urban surface and the atmosphere. Traditional urban canopy models (UCMs) employ tile-based approaches, aggregating aerodynamic drag, radiative trapping, and energy exchange into bulk parameters such as roughness length, displacement height, and plan area density (Lipson et al. 2024). While vertically distributed models (Sützl et al. 2021; Lu et al. 2024) improve momentum and scalar flux partitioning across canopy depths, they assume horizontal homogeneity within grid cells. This assumption breaks down at $O$(1 km) resolutions, as urban landscapes exhibit heterogeneity at neighbourhood scales (about 100-500 m) due to variations in building height, street geometry, and land cover (Stewart and Oke 2012). Such fine-scale variability drives localised microclimates, including urban heat islands, airflow channelling, and pollutant dispersion (Masson et al. 2020), which are poorly captured by current parameterizations.
@@
In the urban canopy region, the flow dynamics are directly influenced by buildings and surface obstacles, governed by drag forces, turbulence generation, and scalar transport. Considering the heterogeneity at neighbourhood scales, moving from traditional 1-D NWP models or UCMs to hectometric models (HMs) can resolve flow structures at $O$(100 m) and, explicitly represent neighbourhood-scale heterogeneities (e.g., building clusters, parks, and street canyons) that drive horizontal variability in momentum sinks and heat fluxes (Sützl et al. 2021; Yu et al. 2023). By resolving these features, for example, in terms of drag, HMs minimise errors introduced by assuming uniform roughness or drag coefficients within a grid cell (Blunn et al. 2022). Recent studies demonstrate that HMs improve predictions of near-surface winds and turbulence kinetic energy in urban areas by up to 30% compared to coarser models (Lean et al. 2024), highlighting their potential to reduce dependency on empirical parameterizations. Moreover, resolving neighbourhood-scale features partially enters the ‘turbulence grey zone’, where subfilter turbulent fluxes coexist with resolved eddies (Honnert et al. 2020). At these scales, models can better represent terrain-driven phenomena (e.g., cold pools in valleys, slope winds) and urban surface-atmosphere interactions (Brun et al. 2017; Smith et al. 2021).
However, challenges persist while pushing current kilometre-scale models toward hectometric-scale models: firstly, computational costs escalate nonlinearly with resolution (Lean et al. 2024), secondly, parameterisations must transition seamlessly from 1-D planetary boundary layer schemes to 3-D turbulence closures (Boutle et al. 2014), and thirdly, observational datasets for validation at sub-kilometre scales remain sparse (Demuzere et al. 2022). To bridge the gap, an intuitive but critical question arises: to what extent do the current kilometre-scale models still hold in hectometric-scale models when resolution increases? Or in other words, how do NWP requirements evolve as resolutions approach the hectometric scale?
@@
The aim of this paper is to present a computationally efficient multi-resolution framework based on the Whitaker (1999) volume-averaging approach and to apply this framework to results from building-resolving large-eddy simulation. This coarse-graining approach allows us to systematically study heterogeneity in the flow statistics. In the context of NWP models, it facilitates analysis of the required drag and turbulence parameterisation for non-building resolving models at different horizontal resolutions. The paper is structured as follows. §2 mainly introduces a multi-resolution spatial averaging framework to grain the coarse field from the original high-resolution one; §3 introduces the simulation details of LES; and §4 presents and discusses the multi-resolution results on drag force and turbulent momentum stress, respectively. The highlights are remarked in the §5.
@@ 2.1 Planar averaging using convolution filters
Figure 1: Definition sketch of domain and filter $\mathcal{A}$. (a) plan view, (b) elevation view. The domain of $\Omega$ is comprised of a fluid subdomain $\Omega_{f}$ (in white), a solid subdomain $\Omega_{b}$ (in grey), and a fluid-solid interface $\partial\Omega_{f}$ (solid black lines). The 3-D normal vectors $\boldsymbol{N}$ point into the fluid domain. The 2-D square filter $\mathcal{A}$ with averaging length $L$ is shown in red.
The volume-averaging framework of Whitaker (1999) as applied to urban environments (Xie and Fuka 2018; Schmid et al. 2019) can be used as a multi-resolution framework by varying the averaging length scale $L$. The theory is well developed but is not straightforward to apply directly to simulation data — analyses are typically performed using averages over the entire domain and not for averages of arbitrary lengthscale $L$. In this section, we present the volume averaging framework in terms of convolution filters. The main advantages of using convolution filters are: 1) that it standardises the methodology, bringing it close to the large-eddy simulation formalism (Pope 2000); and 2) that convolutions can be computed efficiently using fast-Fourier transforms (FFTs). In addition, we formulate the spatial averaging purely in terms of a planar filter. The work in Schmid et al. (2019) is cast in planar form, but the integral representing the effect of the solid boundaries is still presented in the volume-averaging form. Here, we capitalise on recent work by Van Reeuwijk et al. (2021) who derive identities that permit a fully planar formulation. A description of how the volume averaging can be transformed to planar averaging and a derivation of the spatial averaging theorem using convolutions is provided in appendix A. The main results needed for the convolution formalism are presented here.
The domain of interest $\Omega$ is comprised of a subdomain containing the fluid $\Omega_{f}$ and a solid subdomain $\Omega_{b}$ such that $\Omega=\Omega_{f}\cup\Omega_{b}$ (figure 1). The boundary between the fluid and solid phase is denoted $\partial\Omega_{f}$ which has a (3D) normal vector $\boldsymbol{N}$ pointing into the fluid domain. The horizontal, spanwise and vertical coordinates are denoted $x$, $y$ and $z$, respectively. Since averaging will take place over the horizontal directions, the coordinate vector $\boldsymbol{x}$ is denoted as $\boldsymbol{x}=[\boldsymbol{x}_{\perp};z]$, where $\boldsymbol{x}_{\perp}=[x,y]^{T}$. Denoting the filter associated with the area-averaging $\mathcal{A}$, the superficial average (also called comprehensive average; cf. Xie and Fuka 2018) of an arbitrary scalar $\varphi$ is given by
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\varphi(\boldsymbol{y}_{\perp},z)\textnormal{d}\boldsymbol{y}_{\perp}\,.$ (1)
@@
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})=\int_{\Omega(z)}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})I_{f}(\boldsymbol{y}_{\perp},z)\varphi(\boldsymbol{y}_{\perp},z)\textnormal{d}\boldsymbol{y}_{\perp}\,.$ (2)
@@
Formally, there are few restrictions on the filter, except that it satisfies the normalisation condition $\int\mathcal{A}(\boldsymbol{x}_{\perp})\textnormal{d}\boldsymbol{x}_{\perp}=1$. Here, a symmetric square filter is used, defined as
$\mathcal{A}(x,y)=\begin{cases}L^{-2}&\textnormal{if }\left(-\frac{L}{2}<x<\frac{L}{2}\right)\cup\left(-\frac{L}{2}<y<\frac{L}{2}\right)\,,\\ 0&\textnormal{otherwise}\,,\end{cases}$ (3)
@@
In the filter formalism, the area (planar) volume fraction $\varepsilon=A_{f}/L^{2}$, where $A_{f}$ is the fluid surface within the averaging region, is defined as
$\varepsilon(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}.$ (4)
This equation shows the interaction between the filter $\mathcal{A}$ and the integration over the fluid domain $\Omega_{f}$, which may only cover part of the area over which $\mathcal{A}$ filters. The porosity $\varepsilon$ can be used to calculate the intrinsic average $\left\langle{\smash{\varphi}}\right\rangle{{}^{f}}$, indicative of the average value of $\varphi$ inside the fluid phase:
@@ 2.2 Differentiation rules
$\left\langle{\smash{\frac{\partial\varphi}{\partial x_{i}}}}\right\rangle=\frac{\partial\left\langle{\smash{\varphi}}\right\rangle}{\partial x_{i}}-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{N_{i}}{|\boldsymbol{N}_{\perp}|}\,\textnormal{d}s\,.$ (6)
@@
Here, the last term is a horizontal line integral over the fluid-solid boundary $\partial\Omega_{f}$ at given $z$ within the support of $\mathcal{A}$. The negative sign of the line integral is due to the (3D) surface normals $\boldsymbol{N}$ pointing into the domain. The quantity $\boldsymbol{N}_{\perp}=[N_{x},N_{y}]^{T}$ is the component of $\boldsymbol{N}$ in the horizontal $x-y$ plane. The term $|\boldsymbol{N}_{\perp}|$ accounts for the local surface orientation that needs not be aligned with the horizontal direction. For vertical surfaces, $|\boldsymbol{N}_{\perp}|=1$, for angled surfaces $10<|\boldsymbol{N}_{\perp}|<$ 1, and for horizontal surfaces (e.g., flat roofs), $|\boldsymbol{N}_{\perp}|=0$. Superficially it seems the integral is undefined for horizontal surfaces, but it can be shown (Van Reeuwijk et al. 2021) that the integral becomes finite upon integrating over $z$. Physically, one can interpret this as a flat surface providing a finite contribution over an infinitesimal distance in $z$, thus taking the form of a Dirac delta function (see also §3.2). Note that the presence of these apparent singularities is an unavoidable consequence of using a planar average rather than a volumetric average.
The vector form of the spatial averaging theorem (Whitaker 1999, eq. 1.2-23) is given by (see appendix A)
@@
$\left\langle{\smash{\frac{\partial F_{i}}{\partial x_{i}}}}\right\rangle=\frac{\partial\left\langle{\smash{F_{i}}}\right\rangle}{\partial x_{i}}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s\,,$ (7)
where repeated indices imply implicit summation. If the averaging length $L$ is much larger than the characteristic length scale $\ell$ of the urban surface, the average concentration $\left\langle{\smash{\varphi}}\right\rangle$ and horizontal fluxes $\left\langle{\smash{F_{x}}}\right\rangle$ and $\left\langle{\smash{F_{y}}}\right\rangle$ will be independent of $x,y$, implying that the spatial averaging theorems reduce to (Schmid et al. 2019)
$\left\langle{\smash{\frac{\partial\varphi}{\partial x_{i}}}}\right\rangle$ $=\frac{\partial\left\langle{\smash{\varphi}}\right\rangle}{\partial z}-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{N_{i}}{|\boldsymbol{N}_{\perp}|}\,\textnormal{d}s\,,$ (8)
@@
$\left\langle{\smash{\frac{\partial F_{i}}{\partial x_{i}}}}\right\rangle$ $=\frac{\partial\left\langle{\smash{F_{z}}}\right\rangle}{\partial z}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{\perp}|}\,\textnormal{d}s.$ (9)

### v2

# Jingzi Huang

@@ Abstract
A computationally efficient multi-scale planar-averaging framework for urban areas is developed, which enables efficient computation of coarse-grained velocity and scalar fields. We apply the multi-scale framework to a large-eddy simulation of an idealised heterogeneous urban environment of 512 buildings based on a typical London height distribution. We observe that for this geometry, the characteristic urban lengthscale $\ell\approx 50$ m, which is the averaging lengthscale $L$ at which as much variance in the mean flow is resolved as is unresolved. For $L>400$ m, the statistics become approximately homogeneous, suggesting that non-building-resolving numerical weather prediction (NWP) models can be applied without modification at resolutions of 400 m and above for the case under consideration. We derive the multi-scale planar- and Reynolds-averaged momentum equation and show that for neutral cases, NWP models require parameterisation of the distributed drag and the unresolved turbulence and dispersive stress. An a priori analysis reveals that the drag parameterisation from Sutzl et al., 2020, Bound-Layer Meteorol., 178: 225–248 holds reasonably well for resolutions $L$ above 200 m. Below this value, the problem becomes inhomogeneous and the parameterisation works less well. The unresolved stresses are well represented by a $k-\omega$ closure with a value of $\omega=0.4\,\textnormal{s}^{-1}$. However, an even more accurate closure can be derived from the Sutzl drag parameterisation that does not require further turbulence information.
@@ 1 Introduction
The resolution of Numerical Weather Prediction (NWP) models has advanced significantly over recent decades, transitioning from synoptic-scale OO(10 km) grids (Bryan et al. 2003) to convection-permitting OO(1 km) systems (Baldauf et al. 2011; Tang et al. 2013). This progress, driven by computational advancements, enables explicit resolution of critical atmospheric processes such as deep convection and orographic flows, reducing reliance on parameterisations (Lean et al. 2024). Kilometre-scale models are now operational for weather forecasting, regional climate downscaling (Prein et al. 2015; Belušić et al. 2020), and air quality assessments. However, a persistent limitation across NWP frameworks is the unresolved representation of urban morphology-buildings, vegetation, and surface heterogeneities that remain subgrid-scale features, necessitating parameterised approximations.
@@
The urban canopy region is at the lowest level of the atmospheric boundary layer and is characterised by strong heterogeneity and complex interactions between the urban surface and the atmosphere. Traditional urban canopy models (UCMs) aggregate aerodynamic drag, radiative trapping, and energy exchange into bulk parameters such as roughness length, displacement height, and plan area density that connect to the lowest atmospheric level directly (Lipson et al. 2024). Whilst this approach works reasonably well for areas with low-rise buildings, it becomes problematic for areas with tall buildings that extend substantially into the atmospheric boundary layer. Vertically distributed UCMs (Schoetter et al. 2020; Sützl et al. 2021; Lu et al. 2024) improve momentum and scalar flux partitioning by distributing drag, heat over the vertical. However, both approaches assume horizontal homogeneity within grid cells, which breaks down below OO(1 km) resolutions, as urban landscapes exhibit heterogeneity at neighbourhood scales (about 100-500 m) due to variations in building height, street geometry, and land cover (Stewart and Oke 2012). Such fine-scale variability drives localised microclimates, including urban heat islands, airflow channelling, and pollutant dispersion (Masson et al. 2020), which are poorly captured by current parameterisations.
@@
High-resolution simulation solves some of these problems. Recent studies demonstrate that hectometric models (HMs) improve predictions of near-surface winds and turbulence kinetic energy in urban areas by up to 30% compared to coarser models (Lean et al. 2024), highlighting their potential to reduce dependency on empirical parameterisations. Moreover, resolving neighbourhood-scale features partially enters the ‘turbulence grey zone’, where subfilter turbulent fluxes coexist with resolved eddies (Honnert et al. 2020). At these scales, models can better represent terrain-driven phenomena (e.g., cold pools in valleys, slope winds) and urban surface-atmosphere interactions (Brun et al. 2017; Smith et al. 2021).
However, at high horizontal NWP resolutions it will become necessary to relax the UCM assumption of horizontal homogeneity, and to understand how the vertical distribution depends on the horizontal resolution. In order to be able to provide the answer to this question, a multi-resolution framework is required, whereby high-resolution simulations that resolve individual buildings can be coarse-grained to arbitrary NWP resolution, such that it is possible to diagnose a priori what parameterisation is needed to represent the buildings within the NWP model. Coarse-graining is a well-established technique to study, e.g. grey-zone turbulence (Honnert et al. 2020), but there is no established multi-scale framework that can be used for the urban canopy.
Coarse-graining methods apply a spatial averaging, or filtering, operation on a fully resolved field in order to obtain a ‘smoothed’ coarse-resolution field (Wyngaard 2004; Honnert et al. 2011). Spatial averaging has been routinely used for urban canopies, inspired by the formalism presented for vegetation canopies (Raupach and Shaw 1982; Finnigan 2000; Nepf 2012). Within the context of urban canopies, Schmid et al. (2019) provide a spatial-averaging framework based on the volume-averaging framework of Whitaker (1999), which was developed for porous media flows. A unique aspect of NWP models is that near the surface (i.e. in the canopy layer), the horizontal resolution of NWP models is much larger than the vertical resolution, as variations in the vertical tend to be much stronger than in the horizontal. Indeed, typical NWP simulations have a resolution of OO(1 m) in the vertical but OO(1 km) in the horizontal; even in the context of a hectometric model, the cell aspect ratio will be at least 100. This suggests that an appropriate averaging region on the canopy layer can be defined as a thin slab, i.e. reducing a three-dimensional volume averaging to a two-dimensional planar averaging approach, which is the approach used in Schmid et al. (2019) and also for vegetation canopies (Raupach and Shaw 1982). However, the Schmid et al. (2019) framework is not straightforward to apply directly to simulation data, since analyses are typically performed using averages over the entire domain and not for averages of arbitrary length scale $L$. In addition, the approach is computationally inefficient and does not explicitly use a planar averaging region.
@@
The aim of this paper is to present a computationally efficient multi-resolution framework based on the Whitaker (1999); Schmid et al. (2019) spatial averaging approach and to apply this framework to results from building-resolving large-eddy simulation. This coarse-graining approach allows us to systematically study heterogeneity in the flow statistics. In the context of NWP models, it facilitates analysis of the required drag and turbulence parameterisation for non-building resolving models at different horizontal resolutions. The paper is structured as follows. §2 mainly introduces a multi-resolution spatial averaging framework to grain the coarse field from the original high-resolution one; §3 introduces the simulation details of LES; and §4 presents and discusses the multi-resolution results on drag force and turbulent momentum stress, respectively. The highlights are remarked in the §5.
@@ 2 Multi-resolution planar-averaging framework
In this section, we derive a multi-resolution framework in terms of convolution filters. The main advantages of using convolution filters are: 1) that it standardises the methodology, bringing it close to the large-eddy simulation formalism (Pope 2000); and 2) that convolutions can be computed efficiently using Fast-Fourier transforms (FFTs). In addition, we formulate the spatial averaging purely in terms of a planar filter. The work in Schmid et al. (2019) is cast in planar form, but the integral representing the effect of the solid boundaries is still presented in the volume-averaging form. Here, we capitalise on recent work by Van Reeuwijk et al. (2021) who derive identities that permit a fully planar formulation. A description of how the volume averaging can be transformed to planar averaging and a derivation of the spatial averaging theorem using convolutions is provided in Appendix A. The main results needed for the convolution formalism are presented here.
@@ 2.1 Planar averaging using convolution filters
Figure 1: Definition sketch of domain and filter $\mathcal{A}$. (a) plan view, (b) elevation view. The domain of $\Omega$ is comprised of a fluid subdomain $\Omega_{f}$ (in white), a solid subdomain $\Omega_{b}$ (in grey), and a fluid-solid interface $\partial\Omega_{f}$ (solid black lines). The building-surface 3-D normal vectors $\boldsymbol{N}$ point into the fluid domain, and the decomposition is shown in (c). The 2-D square filter $\mathcal{A}$ with averaging length $L$ is shown in red.
The domain of interest $\Omega$ is comprised of a subdomain containing the fluid $\Omega_{f}$ and a solid subdomain $\Omega_{b}$ such that $\Omega=\Omega_{f}\cup\Omega_{b}$ (Fig. 1). The boundary between the fluid and solid phase is denoted $\partial\Omega_{f}$, which has a (3-D) normal vector $\boldsymbol{N}$ pointing into the fluid domain. The horizontal, spanwise and vertical coordinates are denoted $x$, $y$ and $z$, respectively. Since averaging will take place over the streamwise directions, the coordinate vector $\boldsymbol{x}$ is denoted as $\boldsymbol{x}=[\boldsymbol{x}_{h};z]$, where $\boldsymbol{x}_{h}=[x,y]^{T}$. Denoting the filter associated with the area-averaging $\mathcal{A}$, the superficial average (also called comprehensive average, the average normalised by the total area including the solid phase; cf. Xie and Fuka 2018) of an arbitrary scalar $\varphi$ is given by
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})\equiv\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\varphi(\boldsymbol{y}_{h},z)\textnormal{d}\boldsymbol{y}_{h}\,.$ (1)
@@
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})=\int_{\Omega(z)}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})I_{f}(\boldsymbol{y}_{h},z)\varphi(\boldsymbol{y}_{h},z)\textnormal{d}\boldsymbol{y}_{h}\,.$ (2)
@@
Formally, there are few restrictions on the filter, except that it satisfies the normalisation condition $\int\mathcal{A}(\boldsymbol{x}_{h})\textnormal{d}\boldsymbol{x}_{h}=1$. Here, a symmetric square filter is used, defined as
[…]
@@
In the filter formalism, the area (planar) fraction $\varepsilon=A_{f}/L^{2}$, where $A_{f}$ is the fluid surface within the averaging region, is defined as
$\varepsilon(\boldsymbol{x})\equiv\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}.$ (4)
This equation shows the interaction between the filter $\mathcal{A}$ and the integration over the fluid domain $\Omega_{f}$, which may only cover part of the area over which $\mathcal{A}$ filters. The porosity $\varepsilon$ can be used to calculate the intrinsic average $\left\langle{\smash{\varphi}}\right\rangle{{}^{f}}$ (the average normalised by only the area of the variable fluid phase), indicative of the average value of $\varphi$ inside the fluid phase:
@@ 2.2 Differentiation rules
$\left\langle{\smash{\frac{\partial\varphi}{\partial x_{i}}}}\right\rangle=\frac{\partial\left\langle{\smash{\varphi}}\right\rangle}{\partial x_{i}}-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{N_{i}}{|\boldsymbol{N}_{h}|}\,\textnormal{d}s\,.$ (6)
@@
Here, the last term is a Leibniz-like horizontal line integral over the fluid-solid boundary $\partial\Omega_{f}$ at given $z$ within the support of $\mathcal{A}$. The negative sign of the line integral is due to the (3D) surface normals $\boldsymbol{N}$ pointing into the domain (the direction of the 3D vector $\boldsymbol{N}$ is shown by the arrow in Fig. 1c). The quantity $\boldsymbol{N}_{h}=[N_{x},N_{y}]^{T}$ is the component of $\boldsymbol{N}$ in the horizontal xyx-y plane. The term $|\boldsymbol{N}_{h}|$ accounts for the local surface orientation that need not be aligned with the horizontal direction. For vertical surfaces, $|\boldsymbol{N}_{h}|=1$ (e.g., the horizontal vectors on the last building in Fig. 1b), for angled surfaces $0<|\boldsymbol{N}_{h}|<1$ (e.g., the vectors on the middle building in Fig. 1b), and for horizontal surfaces, for example, flat roofs, $|\boldsymbol{N}_{h}|=0$ (the vertical vector on the last building in Fig. 1b). Superficially, it seems the integral is undefined for horizontal surfaces, but it can be shown (Van Reeuwijk et al. 2021) that the integral becomes finite upon integrating over $z$. Physically, one can interpret this as a flat surface providing a finite contribution over an infinitesimal distance in $z$. Note that the presence of these apparent singularities is an unavoidable consequence of using a planar average rather than a volumetric average, and its evaluation is discussed in §3.2. To demonstrate the physical meaning of the line integral, it is useful to apply (6) to the mean pressure $\overline{p}$. The horizontal pressure gradient then becomes
@@
$\left\langle{\smash{\frac{\partial\overline{p}}{\partial x}}}\right\rangle=\frac{\partial\left\langle{\smash{\overline{p}}}\right\rangle}{\partial x}-\oint_{\partial\Omega_{f}}\mathcal{A}\overline{p}\frac{N_{x}}{|\boldsymbol{N}_{h}|}\,\textnormal{d}s\,.$
This demonstrates that the filtering operation produces the form drag in the $x$-direction within the filter support. Note that under the assumption of flow in the positive $x$-direction and cuboids that face the wind head-on, we would have $N_{x}/|\boldsymbol{N}_{h}|=-1$ on windward vertical surfaces, +1 on leeward vertical surfaces and 0 on sidewalls. The top surface of buildings will not contribute to the form drag since $N_{x}=0$, although care must be taken to show that the apparent singularity caused by $|\boldsymbol{N}_{h}|$ = 0 does not cause trouble (see §3.2).
For the planar filter $\mathcal{A}$, the vector form of the spatial averaging theorem (Whitaker 1999, eq. 1.2-23) is given by (see appendix A)
@@
$\left\langle{\smash{\frac{\partial F_{i}}{\partial x_{i}}}}\right\rangle=\frac{\partial\left\langle{\smash{F_{i}}}\right\rangle}{\partial x_{i}}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{h}|}\textnormal{d}s\,$, (7)
where repeated indices imply implicit summation. As an example of the application of this theorem, consider filtering the mean continuity equation $\nabla\cdot\overline{\boldsymbol{u}}=0$. Using (7) with $F_{i}=\overline{u}_{i}$ results in
$\left\langle{\smash{\frac{\partial\overline{u}_{i}}{\partial x_{i}}}}\right\rangle=\frac{\partial\left\langle{\smash{\overline{u}_{i}}}\right\rangle}{\partial x_{i}}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{\overline{u}_{i}N_{i}}{|\boldsymbol{N}_{h}|}\textnormal{d}s=\frac{\partial\left\langle{\smash{\overline{u}_{i}}}\right\rangle}{\partial x_{i}}\,$,
since $\overline{u}_{i}N_{i}=0$ on $\partial\Omega_{f}$ due to the impermeability of the buildings. Therefore, $\nabla\cdot\left\langle{\smash{\overline{\boldsymbol{u}}}}\right\rangle=0$, implying that the filtered velocity field is also divergence-free.
If the averaging length $L$ is much larger than the characteristic length scale $\ell$ of the urban surface, the average concentration $\left\langle{\smash{\varphi}}\right\rangle$ and horizontal fluxes $\left\langle{\smash{F_{x}}}\right\rangle$ and $\left\langle{\smash{F_{y}}}\right\rangle$ will be independent of $x,y$, implying that the spatial averaging theorems reduce to (Schmid et al. 2019)
$\left\langle{\smash{\frac{\partial\varphi}{\partial x_{i}}}}\right\rangle$ $=\frac{\partial\left\langle{\smash{\varphi}}\right\rangle}{\partial z}-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{N_{i}}{|\boldsymbol{N}_{h}|}\,\textnormal{d}s\,$, (8)
$\left\langle{\smash{\frac{\partial F_{i}}{\partial x_{i}}}}\right\rangle$ $=\frac{\partial\left\langle{\smash{F_{z}}}\right\rangle}{\partial z}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{F_{i}N_{i}}{|\boldsymbol{N}_{h}|}\,\textnormal{d}s.$ (9)